# HybridNeRF: Efficient Neural Rendering via Adaptive Volumetric Surfaces

### v1

@@ 4 Experiments
As our goal is high-fidelity view synthesis at VR resolution ($\approx$ 4 megapixels), we primarily evaluate HybridNeRF against the Eyeful Tower dataset [35], which contains high-fidelity scenes designed for walkable VR (Section 4.2). We compare our work to a broader range of methods on additional datasets in Section 4.3. We ablate our design decisions in Section 4.4.
@@ 4.1 Implementation
We train our models in the PyTorch framework [24] and implement our renderer in C++/CUDA. We parameterize unbounded scenes with MERF’s piecewise-linear contraction [27] so that our renderer can efficiently query the occupancy grid via ray-AABB intersection. We train on each scene for 200,000 iterations (100,000 in each training stage) with 12,800 rays per batch using Adam [15] and a learning rate of $2.5\!\times\!10^{-3}$.
@@
Table 2: MipNeRF 360 [2]. Real-time methods are highlighted (best, second-best). Baseline numbers are reported from [27, 37, 5]. MobileNeRF [5] was not evaluated on indoor scenes. Our method performs similarly to SOTA across real-time and offline methods.
@@ 4.2 VR Rendering
We compare HybridNeRF to baselines across the fidelity/speed spectrum. We benchmark several volumetric methods, including (1) iNGP [21], (2) VR-NeRF [35], which extends iNGP’s [21] primitives to better handle HDR reconstruction and provides the best existing results against the Eyeful Tower dataset in terms of visual quality, and (3) MERF [27], a highly optimized method that uses sampling and memory layout optimizations to accelerate rendering. We also compare to VolSDF [36] as a hybrid surface–volume method similar to the first stage of our method. As the original VolSDF implementation uses large MLPs that are unsuitable for real-time rendering, we use an optimized version built on top of iNGP’s acceleration primitives as a fairer comparison. We finally compare against 3D Gaussian splatting [12] as a non-neural approach that represents the current state of the art with regards to quality and rendering speed.
@@ Results
We summarize our results in Table 1 along with qualitative results in Figure 7. VR-NeRF [35] and iNGP [21] render well below real-time frame rates. Our VolSDF implementation, which uses the same primitives as iNGP, is 3$\times$ faster merely from the benefits of using a surface representation (and fewer samples per ray). MERF [27], as a volume representation, relies instead on precomputation to accelerate rendering by explicitly storing diffuse color and density outputs during its baking stage and using only a small MLP to model view-dependent effects. Although it reaches a high frame rate, it provides the least visually appealing results amongst our baselines. 3D Gaussian splatting [12] renders the fastest, but struggles with shadows and lighting changes across the training views and models them as unsightly floaters. Our method is the only to achieve both high quality and real-time frame rates.
@@ 4.3 Additional Comparisons
We compare HybridNeRF to a wide set of baselines on Mip-NeRF 360. We use the same set of baselines as in Section 4.2 for ScanNet++.
@@ Results
We list results in Table 2 and Table 4. Our method performs comparably to the state-of-the-art on Mip-NeRF 360 across both real-time [12] and offline [2] methods. Although ScanNet++ [38] contains fewer lighting inconsistencies across training images than the Eyeful Tower dataset [35], 3D Gaussian splatting still struggles to reconstruct specular surfaces (whiteboards, reflective walls) and backgrounds (Table 4). Our method performs the best overall while maintaining a real-time frame rate.
Table 4: ScanNet++ [38] results. Similar to Table 1, our method is the only to hit VR rate rates along with 3DGS and MERF. HybridNeRF performs best across all quality metrics.
@@
We present results against the Eyeful Tower [35] in Table 3. Spatially adaptive surfaceness is crucial as using a global parameter degrades either speed (when $\beta$ is optimized for quality) or rendering quality (when set for speed). Applying uniform Eikonal loss instead of our distance-adjusted variant degrades quality in unbounded scenes. Omitting the distillation process has a minor impact on quality relative to rendering speed. We note a similar finding when using iNGP [21] primitives instead of CUDA textures, which suggests that introducing hardware acceleration into these widely used primitives is a potential avenue for future research.
@@ 5 Limitations
Storing features in dense 3D grids and triplanes consumes significantly more memory than with hash tables [21]. Training is especially memory-intensive as intermediate activations must be stored for backpropagation along with per-parameter optimizer statistics. Storing features in a hash table during the training phase before “baking” them into explicit textures as in MERF [27] would ameliorate train-time consumption but not inference-time consumption.
@@
Although our training time is much faster than the original NeRF, it is about 2$\times$ slower than iNGP due to the additional backprogation needed for Eikonal regularization (inline with other “fast” surface approaches such as NeuS-facto [41]) and slower than 3D Gaussian Splatting.
@@ 6 Conclusion
We present a hybrid surface–volume representation that combines the best of surface and volume-based rendering into a single model. We achieve state-of-the-art quality across several datasets while maintaining real-time framerates at VR resolutions. Although we push the performance frontier of raymarching approaches, a significant speed gap remains next to splatting-based approaches [12]. Combining the advantages of our surface–volume representation with these methods is a valuable next step.
@@ Appendix A Color Distillation
We distill the MLP used to represent distance from our 256-wide MLP to a 16-wide network during the finetuning stage (Section 3.2). It is possible to further accelerate rendering by similarly distilling the color MLP. We found this to provide a significant boost in rendering speed (from 46 to 60 FPS) at the cost of a minor but statistically significant decrease in rendering quality (see Table 5). We observed qualitatively similar results when decreasing width from 64 to 32 channels with more notable changes in color when decreasing the width to 16 channels (see Figure 9). As our initial results suggest that MLP evaluation remains a significant rendering bottleneck, replacing our scene-wide color MLP with a collection of smaller, location-specific MLPs as suggested by KiloNeRF [26] is potential future work that could boost rendering speed at a smaller cost in quality.
@@ Appendix B Anti-Aliasing
Appendix C ScanNet++
We evaluate 9 scenes from ScanNet++ [38] in Section 4.3 (5fb5d2dbf2, 8b5caf3398, 39f36da05b, 41b00feddb, 56a0ec536c, 98b4ec142f, b20a261fdf, f8f12e4e6b, fe1733741f). We undistort the fisheye DSLR captures to pinhole images using the official dataset toolkit [39] to facilitate comparisons against 3D Gaussian splatting [12] (whose implementation does not support fisheye projection). We use the official validation splits, which consist of entirely novel trajectories that present a more challenging novel-view synthesis problem than the commonly used pattern of holding out every eighth frame [19]. The dataset authors note that their release is still in the beta testing phase, and that the final layout is subject to change. Our testing reflects the dataset as of November 2023.
Figure 9: Color Distillation. Distilling the color MLP to a smaller width during the finetuning stage (Section 3.2) accelerates rendering at the cost of a minor decrease in quality. We observe largely similar results when decreasing the width to 32 channels, and more noticeable changes in color when further decreasing to 16.
@@ Appendix C ScanNet++
Appendix D Societal Impact

### v2

@@ 4 Experiments
As our goal is high-fidelity view synthesis at VR resolution ($\approx$ 4 megapixels), we primarily evaluate HybridNeRF against the Eyeful Tower dataset [38], which contains high-fidelity scenes designed for walkable VR (Sec. 4.2). We compare our work to a broader range of methods on additional datasets in Sec. 4.3. We ablate our design in Sec. 4.4.
@@ 4.1 Implementation
We train our models in the PyTorch framework [26] and implement our renderer in C++/CUDA. We parameterize unbounded scenes with MERF’s piecewise-linear contraction [29] so that our renderer can query the occupancy grid via ray-AABB intersection. We train on each scene for 200,000 iterations (100,000 in each training stage) with 12,800 rays per batch using Adam [17] and a learning rate of $2.5\!\times\!10^{-3}$.
@@
Table 2: MipNeRF 360 [2]. Real-time methods are highlighted (best, second-best, third-best). Baseline numbers as published [29, 40, 6]. MobileNeRF [6] was not evaluated on indoor scenes. Our method performs similar to state-of-the-art real-time and offline methods.
@@ 4.2 VR Rendering
We compare HybridNeRF to baselines across the fidelity/speed spectrum. We benchmark several volumetric methods, including (1) iNGP [23], (2) VR-NeRF [38], which extends iNGP’s [23] primitives to better handle HDR reconstruction, (3) Zip-NeRF [3], an anti-aliasing method that generates high-quality renderings at the cost of speed, and (4) MERF [29], a highly optimized method that uses sampling and memory layout optimizations to accelerate rendering. We also compare to VolSDF [39] as a hybrid surface–volume method similar to the first stage of our method. As the original VolSDF implementation uses large MLPs that are unsuitable for real-time rendering, we use an optimized version built on top of iNGP’s acceleration primitives as a fairer comparison. We also benchmark 3D Gaussian splatting [14] as a non-neural approach that represents the current state of the art with across rendering quality and speed.
@@ Results
We summarize our results in Tab. 1 along with qualitative results in Fig. 7. VR-NeRF [38], iNGP [23], and Zip-NeRF [3] render well below real-time frame rates. Our VolSDF implementation, which uses the same primitives as iNGP, is 3$\times$ faster merely from the benefits of using a surface representation (and fewer samples per ray). MERF [29], as a volume representation, relies instead on precomputation to accelerate rendering by explicitly storing diffuse color and density outputs during its baking stage and using only a small MLP to model view-dependent effects. Although it reaches a high frame rate, it provides the least visually appealing results amongst our baselines. 3D Gaussian splatting [14] renders the fastest, but struggles with shadows and lighting changes across the training views and models them as unsightly floaters. Our method is the only to achieve both high quality and real-time frame rates.
@@ 4.3 Additional Comparisons
We compare HybridNeRF to a wide set of baselines on Mip-NeRF 360. We use the same set of baselines as in Sec. 4.2 for ScanNet++.
@@ Results
We list results in Tab. 2 and Tab. 4. Our method performs comparably to the best on Mip-NeRF 360 across both real-time [14] and offline [2] methods. Although ScanNet++ [41] contains fewer lighting inconsistencies across training images than the Eyeful Tower dataset [38], 3D Gaussian splatting still struggles to reconstruct specular surfaces (whiteboards, reflective walls) and backgrounds (Tab. 4). Our method performs the best amongst real-time methods and comparably to Zip-NeRF [3], while rendering >400$\times$ faster.
Table 4: ScanNet++ [41] results. Similar to Tab. 1, our method is the only to hit VR FPS rates along with 3DGS and MERF. Our quality is near-identical to Zip-NeRF while rendering >400$\times$ faster.
@@
We present results against the Eyeful Tower [38] in Tab. 3. Spatially adaptive surfaceness is crucial as using a global parameter degrades either speed (when $\beta$ is optimized for quality) or rendering quality (when set for speed). Applying uniform Eikonal loss instead of our distance-adjusted variant degrades quality in unbounded scenes. Omitting the distillation process has a minor impact on quality relative to rendering speed. We note a similar finding when using iNGP [23] primitives instead of CUDA textures, which suggests that introducing hardware acceleration into these widely used primitives is a potential avenue for future research.
@@ 5 Limitations
Storing features in dense 3D grids and triplanes consumes significantly more memory than with hash tables [23]. Training is especially memory-intensive as intermediate activations must be stored for backpropagation along with per-parameter optimizer statistics. Storing features in a hash table during the training phase before “baking" them into explicit textures as in MERF [29] would ameliorate training-time consumption but not at inference time.
@@
Although our training time is much faster than the original NeRF, it is about 2$\times$ slower than iNGP due to the additional backprogation needed for Eikonal regularization (in line with other “fast" surface approaches such as NeuS-facto [44]), and slower than 3D Gaussian splatting.
@@ 6 Conclusion
We present a hybrid surface–volume representation that combines the best of surface and volume-based rendering into a single model. We achieve state-of-the-art quality across several datasets while maintaining real-time frame rates at VR resolutions. Although we push the performance frontier of raymarching approaches, a significant speed gap remains next to splatting-based approaches [14]. Combining the advantages of our surface–volume representation with these methods would be a valuable next step.
@@ Appendix A Color Distillation
We distill the MLP used to represent distance from our 256-wide MLP to a 16-wide network during the finetuning stage (Sec. 3.2). It is possible to further accelerate rendering by similarly distilling the color MLP. We found this to provide a significant boost in rendering speed (from 46 to 60 FPS) at the cost of a minor but statistically significant decrease in rendering quality (see Tab. 5). We observed qualitatively similar results when decreasing width from 64 to 32 channels with more notable changes in color when decreasing the width to 16 channels (see Fig. 9). As our initial results suggest that MLP evaluation remains a significant rendering bottleneck, replacing our scene-wide color MLP with a collection of smaller, location-specific MLPs, as suggested by KiloNeRF [28] and SMERF [7], is potential future work that could boost rendering speed at a smaller cost in quality.
@@ Appendix B Anti-Aliasing
Appendix C Model architecture
We render color and distance as follows:
$\boldsymbol{c}(\boldsymbol{x},\boldsymbol{d})$ $=\text{MLP}_{\text{col}}(\Gamma_{\text{col}}(\boldsymbol{x}),\text{SH}(\boldsymbol{d}))$ (9)
$f(\boldsymbol{x})$ $=\text{MLP}_{\text{dist}}(\Gamma_{\text{dist}}(\boldsymbol{x}))\text{,}$ (10)
where $\Gamma_{\text{col}}$ and $\Gamma_{\text{dist}}$ are separate spatial feature encodings:
$\Gamma_{\bullet}(\boldsymbol{x})$ $=\bigoplus_{g\in\{G_{\bullet},T_{\bullet}^{1},T_{\bullet}^{2},T_{\bullet}^{3}\}}\sum_{l=0}^{L_{g}-1}w_{l}\cdot g(\boldsymbol{x},l)\text{.}$ (11)
Here, $L_{g}$ is the number of levels in the 3D grid $G_{\bullet}$ and triplanes $\{T_{\bullet}^{1},T_{\bullet}^{2},T_{\bullet}^{3}\}$, $g(\boldsymbol{x},l)$ is the (interpolated) feature vector at $\boldsymbol{x}$ for level $l$, $w_{l}$ is a per-level dampening weight for anti-aliasing (Eq. 8) and $\bigoplus$’ is concatenation. We encode the direction $\boldsymbol{d}$ through spherical harmonics, $\text{SH}(\boldsymbol{d})$, as an auxiliary input to the color MLP (Eq. 9) that is independent of the feature vector $\Gamma_{\text{col}}(\boldsymbol{x})$.
Figure 9: Color Distillation. Distilling the color MLP to a smaller width during the finetuning stage (Sec. 3.2) accelerates rendering at the cost of a minor decrease in quality. We observe largely similar results when decreasing the width to 32 channels, and more noticeable changes in color when further decreasing to 16.
@@ Appendix C Model architecture
Table 6: Grid feature layout. We measure the effect of using only 3D or triplane features on the Eyeful Tower dataset [38], and note a significant drop in quality when compared to using both.
Pinhole Fisheye Overall
[…]
3D Only 27.10 0.832 0.410 32.17 0.928 0.187 29.41 0.875 0.308
Triplane Only 28.24 0.843 0.312 33.16 0.938 0.150 30.47 0.886 0.238
Both 29.07 0.880 0.268 34.57 0.952 0.115 31.57 0.913 0.198
We use low-resolution 3D grids and high-resolution triplanes, as in previous work (MERF [29]) to obtain the best rendering quality (Tab. 6). We double the grid resolution between levels and therefore have a low-resolution 3D grid with 3 levels ($128^{3}$$512^{3}$) and higher-resolution triplanes with 7 levels ($128^{3}$$8192^{3}$). Naïvely computing Eq. 11 requires $3\!+\!3\!\times\!7=24$ texture fetches (we rely on the CUDA texture API for hardware interpolation and do not need to explicitly query voxel corners/texels). As a render-time optimization, we save pre-summed features $g^{\prime}_{L}$ for each resolution level $L$ (where we store $g^{\prime}_{L}(\boldsymbol{v})=\sum_{l=0}^{L}g(\boldsymbol{v},l)$ for each texel $\boldsymbol{v}$ in $L$), such that Eq. 11 can be rewritten as $\Gamma(\boldsymbol{x})=\bigoplus_{g\in\{G,T^{1},T^{2},T^{3}\}}[w_{L}g^{\prime}_{L}(\boldsymbol{x})+(1-w_{L})g^{\prime}_{L-1}(\boldsymbol{x})]$ for $L\!=\!L(\boldsymbol{x})$ (Eq. 7). Here, $\boldsymbol{v}$ refers to the texel (voxel). Querying two levels requires only $2+3\times 2=8$ texture fetches.
Appendix D Geometric Reconstruction
We evaluate geometric reconstruction on ScanNet++ [41] (which has “ground-truth" laser scan depth only for foreground pixels) in Tab. 7 for the strategies in Fig. 6. Using uniform Eikonal loss in contracted space degrades accuracy (0.419 m error vs 0.219 m for uniform world space and 0.221 m with our distance-adjusted method) and omitting Eikonal loss gives the worst results (0.996 m).
Appendix E ScanNet++
We evaluate 9 scenes from ScanNet++ [41] in Sec. 4.3 (5fb5d2dbf2, 8b5caf3398, 39f36da05b, 41b00feddb, 56a0ec536c, 98b4ec142f, b20a261fdf, f8f12e4e6b, fe1733741f). We undistort the fisheye DSLR captures to pinhole images using the official dataset toolkit [42] to facilitate comparisons against 3D Gaussian splatting [14] (whose implementation does not support fisheye projection). We use the official validation splits, which consist of entirely novel trajectories that present a more challenging novel-view synthesis problem than the commonly used pattern of holding out every eighth frame [21]. The dataset authors note that their release is still in the beta testing phase, and that the final layout is subject to change. Our testing reflects the dataset as of November 2023.
Table 7: Depth error on ScanNet++ [41]. Our distance-adjusted Eikonal loss degrades geometric reconstruction less than other alternatives used to render unbounded scenes.
[…]
Uniform Eikonal loss (world space) 0.219 8.56
Uniform Eikonal loss (contracted space) 0.419 16.11
No Eikonal loss 0.996 29.93
Distance-adjusted Eikonal loss (ours) 0.221 11.13
Appendix F Societal Impact